Lilly2010:

Find the relationships of rate of change (slope) and initial value when x=0 (y intercept) between x and y for the following table of values of x, the independent variable and y, the dependent variable. Also find the equation of the line in slope intercept form, y=mx+b. Please show and explain your work

Lilly2010:

Lilly2010:

🙏

unsolvedlife:

The rate of change is -2 and the initial value when x=0 (y intercept) is -1. An equation of this line in slope-intercept form is y = -2x - 1.

563blackghost:

To identify the rate of change you can identify it by the following formula: \(m=\frac{y_{2}-y_{1}}{x_{2}-x_{1}}\) You can input two points into the equation let's use the first two points. \(m=\frac{3-5}{(-1)-(-2)}\) Simplify to identify the slope (rate of change).

unsolvedlife:

@563blackghost wrote:
To identify the rate of change you can identify it by the following formula: \(m=\frac{y_{2}-y_{1}}{x_{2}-x_{1}}\) You can input two points into the equation let's use the first two points. \(m=\frac{3-5}{(-1)-(-2)}\) Simplify to identify the slope (rate of change).
mb about posting the answer

Lilly2010:

Thank you all so much for your help! it is appreciated.